OpenStudy (blackfireskull):

someone please help me .-. Evaluate 6 - 2(-1) + | -5 | = 1 9 13

OpenStudy (blackfireskull):

@Goose. i think it's 13. but not 100% sure .-.

Nnesha (nnesha):

i just want to ask one question \[\left| -5 \right| = ?\]

OpenStudy (alesia/bailey):

I got 13 :P

OpenStudy (blackfireskull):

5. because when you see || it means the absolute Value of the number in between those lines. @Nnesha

OpenStudy (blackfireskull):

and that's exactly what i got as well @alesia/bailey

Nnesha (nnesha):

yes and you will get positive number always!! good job!

OpenStudy (goose.):

May I ask how you got 13? I got 9...

OpenStudy (blackfireskull):

6-2(-1)+|-5|= 6-2(-1)+5

OpenStudy (goose.):

Oh wait, never mind I see what I did, haha.

OpenStudy (blackfireskull):

lmao wooooow Goose XD

OpenStudy (goose.):

After distributing the -2 I forgot to drop my '-' sign.

Nnesha (nnesha):

and remember the PEMDAS rule multiply /divide first and then subtract /add

OpenStudy (goose.):

Multiplication or Division in the order of which it appears, Addition or Subtraction order of which it appears.